OpenStudy (anonymous):

What is the next fraction in this sequence? Simplify your answer. 4/69 , 5/23, 26/69, 37/69, ... 15/23 44/69 52/69 16/23

OpenStudy (hoblos):

48/69 = 16/23

OpenStudy (anonymous):

I don't get it

OpenStudy (hoblos):

you are adding 11/69 each time 37/69 + 11/69 = 48/69=16/23

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

welcome -_-